# Few-Shot Classification of Interactive Activities of Daily Living (InteractADL)

### v1

@@ 1 Introduction
Activities of Daily Living (ADLs) are defined as a set of basic, essential, and routine tasks that most healthy people can perform without any assistance [Katz(1983)]. Recognizing ADLs is an essential component in several applications, from smart homes [Debes et al.(2016)Debes, Merentitis, Sukhanov, Niessen, Frangiadakis, and Bauer] and assistive robots [Lauretti et al.(2017)Lauretti, Cordella, Guglielmelli, and Zollo] to automated healthcare assessment [Desai et al.(2004)Desai, Grossberg, and Sheth] applications. Despite these essential use cases, there are few benchmarks and methods for understanding human ADLs at home, especially where there is interaction with other humans involved. Prior work has focused on ADLs at home with a single person present, such as Toyota Smart Home [Dai et al.(2022)Dai, Das, Sharma, Minciullo, Garattoni, Bremond, and Francesca] and Home Action Genome (HOMAGE) [Rai et al.(2021)Rai, Chen, Ji, Desai, Kozuka, Ishizaka, Adeli, and Niebles], or has presented benchmarks for ego-view scenes in specific rooms in the house, e.g., in the kitchen as in EPIC-Kitchens [Damen et al.(2018)Damen, Doughty, Farinella, Fidler, Furnari, Kazakos, Moltisanti, Munro, Perrett, Price, et al.].
@@ 7 Conclusion
In this paper, we introduced a new dataset, InteractADL, which includes multi-view data of multiple humans performing ADLs at home. To tackle the few-shot activity recognition problem posed by this dataset, we introduce a novel method, Name Tuning, for learning model-specific class names that expands upon existing prompt tuning strategies. Through this method, we also achieve a new state-of-the-art in three different few-shot activity recognition benchmarks (MOMA-Activities, MOMA-Subactivities, and Kinetics). We demonstrate the challenges for adapting pre-trained VLMs for our fine-grained video dataset, despite their impressive zero-shot performance on relatively high-level datasets. For transparency, reproducibility, and to accelerate research in this area, we release our code at the following URL.

### v2

@@ 1 Introduction
Activities of Daily Living (ADLs) are defined as a set of basic, essential, and routine tasks that most healthy people can perform without any assistance [Katz(1983)]. Recognizing ADLs is an essential component in several applications, from smart homes [Debes et al.(2016)Debes, Merentitis, Sukhanov, Niessen, Frangiadakis, and Bauer] and assistive robots [Lauretti et al.(2017)Lauretti, Cordella, Guglielmelli, and Zollo] to automated healthcare assessment [Desai et al.(2004)Desai, Grossberg, and Sheth] applications. However, there are few benchmarks and methods for understanding human ADLs at home, especially where there is interaction with other humans. Prior work has focused on ADLs at home with a single person present, such as Toyota Smart Home [Dai et al.(2022)Dai, Das, Sharma, Minciullo, Garattoni, Bremond, and Francesca] and Home Action Genome (HOMAGE) [Rai et al.(2021)Rai, Chen, Ji, Desai, Kozuka, Ishizaka, Adeli, and Niebles], or benchmark ego-view scenes in specific rooms in the house, e.g., in the kitchen as in EPIC-Kitchens [Damen et al.(2018)Damen, Doughty, Farinella, Fidler, Furnari, Kazakos, Moltisanti, Munro, Perrett, Price, et al.].
@@ 7 Conclusion
In this paper, we introduced a new dataset, InteractADL, which includes multi-view data of multiple humans performing ADLs at home. To tackle the few-shot activity recognition problem posed by this dataset, we introduce a novel method, Name Tuning, for learning model-specific class names that expands upon existing prompt tuning strategies. Through this method, we also achieve a new state-of-the-art in three different few-shot activity recognition benchmarks (MOMA-Activities, MOMA-Subactivities, and Kinetics). We demonstrate the challenges for adapting pre-trained VLMs for our fine-grained video dataset, despite their impressive zero-shot performance on relatively high-level datasets. For transparency, reproducibility, and to accelerate research in this area, we release our code at https://github.com/zanedurante/vlm_benchmark.
Acknowledgments
Funds to support this AITC study were provided by the Johns Hopkins University AITC under award number P30AG073104. This was also partially supported by Panasonic and the Jaswa Innovator Award. This article solely reflects the opinions and conclusions of its authors and not any entity associated with Johns Hopkins University AITC or Panasonic.